OpenStudy (anonymous):

Please explain the equation of the circle when the endpoints are (-3,4) and (6,1). (x-a)^2+(y-b)^2=r^2....Help I can not solve

OpenStudy (phi):

I think the points they give you are on a diameter. The center of the circle is the point in between (-3,4) and (6,1). can you find the mid point ?

OpenStudy (phi):

the mid point will have the average x value and the average y value. can you find the average of the x's -3 and 6 ?

OpenStudy (anonymous):

m=-3+6/2 ?

OpenStudy (phi):

the x value of the center point (exactly between -3 and +6) is (-3+6)/2 you get 3/2 or 1.5 can you find the y value of the center point ? the average of the end point's y's

OpenStudy (anonymous):

yes, I will try

OpenStudy (anonymous):

(4+1)/2 I get 5/2 ?

OpenStudy (phi):

looks good, the center of the circle is (3/2 , 5/2) so far we have (x- 3/2 )^2 + (y- 5/2)^2 = r^2 we need the distance from the point (3/2 , 5/2) to one of the end points, for example (6,1) do you know how to find the distance between 2 points ?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

r^2 = (3/2 - 6)^2 + (5/2 - 1)^2

OpenStudy (anonymous):

thank you so much for thanking the time to help me! :)

OpenStudy (phi):

Can you find r^2 ?

OpenStudy (phi):

Here is a graph of your circle

OpenStudy (anonymous):

the r=22.5

OpenStudy (phi):

r^2 is 22.5 (r would be sqrt(22.5) the final equation is (x- 3/2 )^2 + (y- 5/2)^2 = 22.5

OpenStudy (anonymous):

You are so Awesome!! Thank you Much..

OpenStudy (anonymous):

phi, I understand better because of you. once again take you.